OpenStudy (anonymous):

sqrt{x (x-2)}-x =-10?

OpenStudy (anonymous):

x (x-2)>0 sqrt{x (x-2)}=x-10 /take the square x(x-2)=x^2-20x+100 18x=100 x=100/18=50/9 Check sqrt50/9*(32/9)=40/9 40/9-50/9=-10/9 Did something wrong :-)

OpenStudy (anonymous):

Cannot see the mistake, too tired I guess

OpenStudy (anonymous):

I guess it means that it has no solutions

OpenStudy (wasiqss):

andras u did it right

OpenStudy (anonymous):

thanks a lot